2. Assume that a researcher randomly selects 14 newborn babies and counts the number of (I point) girls, x. The probabilities corresponding to the 14 posible values ofx are summarized in the given table. Answer the questions using the table. Probabilities of Girls xigirls) Px) xgirls) P) K(girls) Px) 004 5 0.122 10 0.061 1 0.001 6 0.183 11 0.022 2 0.006 7 0.209 12 0.006 3 0.022 8 0.183 13 0.001 4 0.0619 0.122 14 0 Find the combined probability of selecting 5, 6, 7, 8, or 9 girls. 0.122 0.610 0.819 00.989
Answer #1

Here we want to find the combined probability of selecting 5, 6, 7, 8, or 9 girls.

Mathematically, P( X >= 5 ) = P( X = 5) + P( X = 6) + P( X = 7) + P( X = 8) + P( X = 9)

From the table, we have

P( X >= 5 ) = 0.122 + 0.183 + 0.209 + 0.183 + 0.122 = 0.819 ( This is the final answer).